OpenStudy (anonymous):

use mathematical induction to prove the statement is true for all positive integers n the integer n^3+2n is divisible by 3 for every positive integer n

OpenStudy (anonymous):

ok, so to prove something using induction whats the first step you need to do?

OpenStudy (anonymous):

theres two steps

OpenStudy (anonymous):

i am honestly not sure..

OpenStudy (anonymous):

ok, the first step is to prove that n(1) is true

OpenStudy (anonymous):

i remember a bit of that it's called your "base" i think. but what do i prove this equation against without it being equal to something?

OpenStudy (anonymous):

this is more of a theorem youre proving

OpenStudy (anonymous):

1^3+2(1)=?. the problems i have done like this have another equation in the "?" area

OpenStudy (anonymous):

so for the first step n(1) = (1)^3 +2(1)= 1+2 = 3, so 3is divisable by 3

OpenStudy (anonymous):

oh i see! yes it is

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

see, its not equal to a number, they want you to show that the equation is equal to a number divisable by 3

OpenStudy (anonymous):

so now that you showed its divisable by 1, you need to show step 2 which is....?

OpenStudy (anonymous):

another fabulous question... sorry i don't know

OpenStudy (anonymous):

lol, well whats the question asking?

OpenStudy (anonymous):

i need to prove the statement true for all positive integers, but i'm going to assume that I shouldn't prove individually each number that is not negative.

OpenStudy (anonymous):

well, you are going to prove each number. youre just not going to enter 2, 3, 4, 5,......into the statement.....so now the question is, do you see a pattern?

OpenStudy (anonymous):

n(2) is 12 and n(3) is 33 so i can tell that they all seem divisible by 3,

OpenStudy (anonymous):

if we showed it is true for 1, and we need to prove its true for EVERY natural number what you need to do is plug in "n+1" in to the statement

OpenStudy (anonymous):

that will give us every natural number!

OpenStudy (anonymous):

would i plug that in where all the "n" is? for example: (n+1)^3+2(n+1)

OpenStudy (anonymous):

EXACTLY!! preform the binomial expansion, combine like terms and thats pretty much it. i gotta run to class now, good luck with any other problems!

OpenStudy (anonymous):

thanks!